Solve the following if 11=20 12=31 13 = 42 14 =??? Choose the correct answer a 55 b 53 c58 d 43
Answer by rothauserc(4718)

the answer is b. 53
here is the rule, given a1a2
a1 + a2 is the tens position and a2 - a1 is the units position, therefore
14 = 1 + 4 = 5 in tens position and 4 -1 = 3 in the units position, therefore
14 = 53